NOAA GOES SATOPS Morning Report: Tuesday, February 15, 2005

Status Report From: National Oceanic and Atmospheric Administration
Posted: Tuesday, February 15, 2005

DAY 46


Secure voice returned to service @ 045/1530Z. < OPS 54 rev 06861 at POGO-C 08:38z Lost. A Category I data loss declared by USER at 08:59z. Smooth data loss was 2175 S/C seconds, 74 playback seconds. During support USER equipment failed to ingest data. Data was not reshipped from the site within Sprint data timelines after fade of support. Message left on the voice mail of the ADO Watch Officer concerning data loss.

All other DMSP operations were nominal over the past 24 hours.


< N17 rev 13746 at Fairbanks 0020z : STX-1 power flagged red high limit (NSTX1PWR-RH, V=98) Renee Smith-Dearring notified and came in to evaluate STX-1 power. M. Fitzmaurice also contacted with J. Devine. SOCC engineers will continue to monitor the STX-1 situation. STX-1 is used for playbacks. No data loss because of the anomaly and SOOC has implemented a backup plan using STX-4 to ensure no data loss will occur.

All other POES operations were nominal over the past 24 hours.


GOES spacecraft have started battery reconditioning in preparation for the upcoming eclipse periods.

< GOES-9 045/1433Z - Momentum wheel 1 current spiked red high with a value of 1.422

All other GOES operations were nominal over the past 24 hours.

Questions or comments to

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter